Daniel Berwick Evans. Outline: (Roughly)

(1) Classical field theory (Euler-Lagrange etc.)

(2) Dirac and Klein-Gordon equations. (Classical.)

(3) Yang-Mills stuff. (U(1) is Maxwell.)

(4) Perturbative field theories, Feynman diagrams etc.

(n) Get physicists to give talks. Resources: QFT and Strings for Mathematicians. Zee, QFT in a Nutshell. Peskin and Shroeder, QFT. Banks, Modern QFT.